OpenStudy (mendicant_bias):

Another basic Differential Equation, problem prompt posted below shortly.

OpenStudy (mendicant_bias):

\[\text{Show that the differential equation is exact and then solve the exact equation.}\] \[[\tan(x) − \sin(x)\sin(y)]dx + [\cos(x)\cos(y)]dy = 0\]

OpenStudy (mendicant_bias):

(Oh, lol, nvm)

OpenStudy (anonymous):

So what's troubling ya about this one? :)

OpenStudy (mendicant_bias):

Nothing yet, just want to put it up in case I have an issue, that way I'll be able to possibly get help quick. I just worry about not figuring it out and nobody being around to potentially help. Working through it momentarily.

OpenStudy (anonymous):

Alrighty, np.

OpenStudy (mendicant_bias):

Alright, this looks like an exact equation where the current body of the problem is an exact differential of some function; we assume that function to be continuous/differentiable at its first partials, so its second partials should be equal to one another, e.g.\[f_{xy}=f_{yx}\]

OpenStudy (mendicant_bias):

Or, more familiar, \[N_{x}-M_{y}=0\]

OpenStudy (mendicant_bias):

\[N_{x}=-\sin(x)\cos(y); \ \ \ M_{y}=-\sin(x)\cos(y).\]

OpenStudy (mendicant_bias):

Thus, the equation is exact. Alright, here's where I start getting hazy on remembering what to do, lol, just one second...

OpenStudy (mendicant_bias):

Oh, okay, this is the integrate/take derivative of thing repeatedly, one sec.

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}N_{x}\ dx=\int\limits_{}^{}-[\sin(x)\cos(y)]dx;\]

OpenStudy (mendicant_bias):

I'm allowed to factor out the cos(y), I think, in this situation?

OpenStudy (anonymous):

Wait, you tried to integrate cos(x)cos(y), right? Then you need to integrate that with respect to y.

OpenStudy (anonymous):

Hm...maybe youre using a method different than Im used to :)

OpenStudy (mendicant_bias):

IDK, lemme check I'm doing things right, lul, don't want to go through all the work to find out I fundamentally was doing things wrong in the first place.

OpenStudy (anonymous):

Unless theres a way of doing this I'm unaware of, you need to choose to integrate \(tan(x) - sin(x)sin(y)\) with respect to x or \(cos(x)cos(y)\) with respect to y

OpenStudy (mendicant_bias):

Yeah, I'm going backwards, I even integrated the wrong thing to begin with, lol. Yeah, you're right.

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}[\tan(x)-\sin(x)\sin(y)]dx=\int\limits_{}^{}\tan(x)dx-\int\limits_{}^{}\sin(x)\sin(y)dx\]

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\tan(x)=\ln|\sec(x)|+c, \ \text{or:} -\ln|\cos(x)|+c\]

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\sin(x)\sin(y)dx=\sin(y)\int\limits_{}^{}\sin(x)dx=\sin(y)[-\cos(x)]+c\]

OpenStudy (mendicant_bias):

Wait a sec, both of those should have a function w.r.t. y

OpenStudy (anonymous):

Except its not quite + C this time.

OpenStudy (anonymous):

Yeah, haha

OpenStudy (mendicant_bias):

\[\ln|\sec(x)|-\sin(y)\cos(x)+g(y)=f(x,y)\] (not sure about what I set it equal to, one minute)

OpenStudy (mendicant_bias):

Now I take the partial with respect to y and set it equal to the second term from the original problem IIRC.

OpenStudy (anonymous):

Yes, in order to solve for g'(y) and eventually g(y) itself.

OpenStudy (mendicant_bias):

\[\frac{\partial}{\partial y}\bigg(\ln|\sec(x)|-\sin(y)\cos(x)+g(y)\bigg)=-\cos(y)\cos(x)+g'(y)\]

OpenStudy (mendicant_bias):

\[-\cos(y)\cos(x)+g'(y)=N=\cos(x)\cos(y).\]So g(y) is some arbitrary constant c, but somewhere in there, I clearly dropped a negative sign.

OpenStudy (anonymous):

Yeah, in the integration the sign shouldve changed.

OpenStudy (anonymous):

Youforgot to write the negative in the 2nd integral, so even though you integrated it correctly, you had the wrong integral to start with.

OpenStudy (mendicant_bias):

Oh, yeah, I see where.

OpenStudy (anonymous):

Yay for g(y) being 0? Lol.

OpenStudy (mendicant_bias):

Alright, so at least the negative didn't affect the overall integral. Yeah, makes things easy but also feels kinda lame/trivial, heh

OpenStudy (mendicant_bias):

Alright, now that I know that, I integrate the second part with respect to y and get some terms plus an unknown function of x, I take the partial of it with respect to x, and so on and so forth. One moment.

OpenStudy (anonymous):

No, youre done, lol.

OpenStudy (mendicant_bias):

Wait wat lol

OpenStudy (anonymous):

So you integrated with respect to x and got: \(ln|secx| + sin(y)cos(x) + g(y)\) Once we found that g(y) = 0, our answer simply is \(ln|secx| + sin(y)cos(x) = c\) Thats all you need really.

OpenStudy (mendicant_bias):

Alright, just gimme a sec. Trying to make sure I get this.

OpenStudy (anonymous):

If you do it the other way, youll get the same result. You can integrate the 2nd part with respect to y and do the partial blah blah with respect to x, but it's simply re-doing your work to get the same answer. Typically, you choose which one is easier to ingrate and go from there, doing the process you did to get an answer. It doesnt need to be done twice.

OpenStudy (mendicant_bias):

Yeah, I got it, makes sense now.

OpenStudy (mendicant_bias):

Now I just need to put it either in the form of an implicit or explicit solution, like-oh, you did that, too, lol. Gonna see if I can make it explicit, one sec.

OpenStudy (anonymous):

Leave it. The final solution is a function equal to c, that is how it should look.

OpenStudy (mendicant_bias):

Alright, cool. Thanks you! Going to go through a couple more problems momentarily.

OpenStudy (anonymous):

No problem :)

OpenStudy (mendicant_bias):

Alright, nevermind, going to go eat dinner really quick, well, wait, maybe; how long do you think you're going to be on? If you're planning on getting out of soon, I'll just stay here and get them done now while there are people around that can help.

OpenStudy (anonymous):

Um, another 2-3 hours. Until the starbucks im at closes >.>

OpenStudy (mendicant_bias):

Hah. Alright, I'll be back in about 30-45 minutes, just need to go to McDonald's or some place not crazy pricey before they close, I'll bbl.

OpenStudy (anonymous):

Kk.